OpenStudy (lord2clash):

What is the distance between the points (3, 7) and (15, 16) on a coordinate plane?

pooja195 (pooja195):

Use the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

pooja195 (pooja195):

Can you tell me what x1=? x2=? y1=? y2=?

OpenStudy (lord2clash):

x = (3, 16) and y = (15, 7)

pooja195 (pooja195):

Good \[\huge d=\sqrt{(16-3)^2+(15-7)^2}\] 16-3=? 15-7=?

OpenStudy (lord2clash):

16-3 = 13 15-7 = 8

pooja195 (pooja195):

13^2=? 8^2=?

OpenStudy (lord2clash):

169 64 (thats what my calculator gave me)

pooja195 (pooja195):

169+64=?

OpenStudy (lord2clash):

233

pooja195 (pooja195):

\[\huge d=\sqrt{233}\]

pooja195 (pooja195):

Well done u found the answer in decimal form it would be 15.27

OpenStudy (lord2clash):

if you round up yes. thank you pooja cx

pooja195 (pooja195):

yw ^_^